1. A manufacturer of processing chips knows that 2% of its chips are defective in some way. Suppose an inspector randomly selects 4 chips for an inspection. Assuming the chips are independent, what is the probability that at least one of the selected chips is defective?

2. A box contains 4 red marbles, 3 green marbles, and 2 blue marbles. One marble is removed from the box and it is not replaced. Another marble is drawn from the box.

What is the probability that the first marble is blue, and the second marble is green?
Answer #1

1)

probability that at least one of the selected chips is defective =1-P(none is defective)=1-(0.98)4 =0.077632

2)

probability that the first marble is blue, and the second marble is green =(2/9)*(3/8)=1/12